# HCOA*: Hierarchical Class-ordered A* for Navigation in Semantic Environments

### v1

@@ Abstract
This paper addresses the problem of robot navigation in mixed geometric and semantic 3D environments. Given a hierarchical representation of the environment, the objective is to navigate from a start position to a goal while minimizing the computational cost. We introduce Hierarchical Class-ordered A* (HCOA*), an algorithm that leverages the environmental hierarchy for efficient path-planning in semantic graphs, significantly reducing computational effort. We use a total order over the semantic classes and prove theoretical performance guarantees for the algorithm. We propose two approaches for higher-layer node classification based on the node semantics of the lowest layer: a Graph Neural Network-based method and a Majority-Class method. We evaluate our approach through simulations on a 3D Scene Graph (3DSG), comparing it to the state-of-the-art and assessing its performance against our classification approaches. Results show that HCOA* can find the optimal path while reducing the number of expanded nodes by 25% and achieving a 16% reduction in computational time on the uHumans2 3DSG dataset.
@@ I Introduction
As robotic sensing technologies advance, enabling robots to perceive vast and diverse information, two fundamental questions arise: What information from this extensive data stream is most important for a given task, and how can the robot effectively utilize this information for decision-making? Hierarchical semantic environment representations, such as 3D Scene Graphs (3DSG) [1, 2, 3], provide rich and structured abstractions that mirror human way of thinking, facilitating the selection and organization of information.
@@
Previous research in hierarchical path-planning has primarily addressed the first question [4, 5, 6, 7]. In [4] the authors introduce Hierarchical Path-Finding A*, a hierarchical A* variant for grid-based maps. Their approach partitions the map into clusters with designated entrance points, which are used for high-level path-planning. The authors in [5] propose a hierarchical graph search algorithm for graphs with edge weights represented as intervals. Similarly, in [7] the authors propose two hierarchical path-planning methods: path-making and alternating search direction.
Semantic path-planning has focused on the second question by incorporating semantics into the decision-making process. In [8] the authors introduce a weighted function that combines the edge cost and the semantic class to determine the optimal path, while considering the total order of the semantic classes. However, computing this weighted function requires global graph properties, which can be computationally demanding. To address this limitation, [9] and [10] propose Class-ordered A*/LPA*, two extensions of traditional A* [11] that efficiently incorporate semantics, based on a total order over the available classes. Building on the approaches of [5] and [9], in this paper, we develop a hierarchical class-ordered path-planning algorithm, enhancing scalability to larger and more complex environments.
@@
Navigation within hierarchical semantic environments has also been explored in prior work. In [12], the authors introduce D-Lite, a set of two greedy, navigation-oriented compression algorithms for 3DSGs. These algorithms compress the graph based on the shortest path in the lowest layer, and the compressed representation is used for robot navigation. In [13], the authors address Task and Motion Planning in 3DSGs using a three-level hierarchical planner consisting of a discrete task planner, a navigation planner for path-planning, and a low-level planner. Their approach focuses on optimizing task planning to reduce computational cost, while path-planning operates on the unpruned 3DSG.
@@
Main Contributions: Our approach integrates task semantics directly into the path-planning process within a unified algorithm. By leveraging the hierarchical structure of the environment, we significantly reduce computational resource demands while ensuring efficient navigation. Our key contributions are as follows:
@@
    We propose two methods for node classification on the higher layers of the hierarchy based on underlying semantic classes: a GNN-based method and a majority-class method.
@@
    We validate our approach on a publicly available 3DSG dataset, demonstrating its effectiveness in real-world scenarios.
@@ II Preliminaries
We model the environment as a hierarchical semantic graph (HSG). Specifically, let $G=(V,E,\mathcal{K})$ be a graph with $n$ layers, where $V$ is the set of nodes, $E\subseteq V\times V$ is the set of edges, and $\mathcal{K}={1,\ldots K}$ represents a set of semantic classes ordered by decreasing priority. We denote the set of layers as $L$, where $\ell=0$ is the lowest layer and $\ell=n$ is the highest layer (root). Each layer $\ell$ forms a connected weighted subgraph $G^{l}=(V^{\ell},E^{\ell},\mathcal{K})$, with an associated weight function $w:E^{\ell}\to\mathbb{R}^{+}$. We assume that each node in layer $\ell-1$, for $\ell=1,\ldots n$, is connected to exactly one node in layer $\ell$, which we refer to as its parent node. More generally, we define an ancestor as a node’s parent or any higher-layer predecessor in the hierarchy. We define the projection function $p:V\times L\to V$ that maps each node to its corresponding ancestor in layer $\ell$. An example of an HSG is the 3D Scene Graph (3DSG) shown in Figure 1.
Each node in $\ell=0$ is assigned a semantic class based on perception data using the function $\phi_{V}^{0}:V^{0}\to\mathcal{K}$. We extend the labeling function $\phi_{V}^{0}$ to layers $\ell\neq 0$. Details on the computation of $\phi_{V}^{\ell}$ for $\ell\neq 0$ will be given in Section V. Additionally, we introduce the function $\phi_{E}^{\ell}:E^{\ell}\to\mathcal{K}$, which assigns semantic classes to the edges of $G$. Specifically, we define the edge classification function for an edge $e=(v,u)$ as $\phi_{E}^{\ell}(e)=\max(\phi_{V}^{\ell}(v),\phi_{V}^{\ell}(u))$, thereby overestimating the edge class.
Let $\Pi(v_{s}^{\ell},v_{g}^{\ell})$ denote the set of all acyclic paths in $G^{\ell}$ from $v_{s}^{\ell}\in V^{\ell}$ to $v_{g}^{\ell}\in V^{\ell}$, and let $\Pi_{k}$ be the subset of $\Pi$ in which the least favorable (i.e., highest) edge class is exactly $k$. Formally,
$\Pi_{k}=\{\pi:N(\pi,k)>0\text{ and }N(\pi,k^{\prime})=0,\forall k^{\prime}>k\},$ (1)
where $N(\pi,k)=|\{e\in\pi:\phi_{E}^{\ell}(e)=k\}|$ is the number of edges of class $k$ in path $\pi$. We further define $\Pi_{k}^{i}\subseteq\Pi_{k}$ consisting of paths that contain exactly $i$ edges of class $k$, that is $\Pi_{k}^{i}=\{\pi\in\Pi_{k}:N(\pi,k)=i\}$. To enable a consistent comparison of paths across different classes, we impose a total order such that $k<\ell\Rightarrow\Pi_{k}^{i}\prec\Pi_{\ell}^{j}$ for all $\text{for all }i,j$ and $i<j\Rightarrow\Pi_{k}^{i}\prec\Pi_{k}^{j}$. This order ensures that any two paths with the same start and goal nodes can be compared.
@@
4:     $g(v^{\ell})\leftarrow\infty;\hskip 5.0pt\theta(v^{\ell})\leftarrow\infty\cdot\mathbf{1}_{K},\quad\forall v^{\ell}\in G\setminus\{v_{s}^{\ell}\}$
[…]
6:     $Q\leftarrow\{v_{s}^{\ell}\}$triangleright Initialize priority queue
@@
26:     for all $\ell^{\prime}<\ell$ do
27:         $G^{\ell^{\prime}}\leftarrow\{v\in G^{\ell^{\prime}}:p(v,\ell)\in\pi^{\ell}\}$
@@ III Problem Formulation
Consider a mobile robot tasked with navigating a complex 3D environment where certain regions have lower traversal priority. The robot is provided with a HSG of the environment (e.g., 3DSG), as outlined in Section II, where the semantic classes of the nodes in layer $\ell=0$ are assigned based on perception and tailored to the specific task (e.g. avoid certain objects). Let $v_{s}\in V^{0}$ and $v_{g}\in V^{0}$ denote the robot’s starting and goal nodes, respectively. The objective is to determine the shortest path within layer $\ell=0$ while minimizing traversal through the least favorable edges. Formally,
$\pi^{*}(v_{s},v_{g})=\operatorname*{arg\,min}\limits_{\pi\in\Pi^{*}(v_{s},v_{g})}{\sum\limits_{e\in\pi}w(e)},$ (2a)
@@ III-A Problem Statement
We propose a hierarchical semantic path-planning algorithm for efficient robot navigation in large-scale environments. The algorithm operates top-down across the layers of the HSG, iteratively computing the optimal semantic path at each layer while pruning nodes that are not included in the selected path. Additionally, we introduce two methods for node classification to predict the semantic classes of higher-layer nodes: a Graph Neural Network (GNN)-based approach and a majority-class method.
@@ IV-A Hierarchical Class-ordered A*
We introduce Hierarchical Class-ordered A* (HCOA*), a hierarchical semantic path-planning algorithm for HSGs. The algorithm begins by finding a path in layer $\ell={n-1}$ and then refining it by recursively applying the same process at the lower levels of the hierarchy. By structuring the search hierarchically, the algorithm performs multiple searches on smaller subgraphs rather than executing a computationally expensive search over $G^{0}$. At each layer, the algorithm utilizes Class-ordered A* (COA*) [9], which finds the shortest path while minimizing the number of least favorable edges through lexicographic comparison.
HCOA* is presented in Algorithm 1. We use the function $h:V\to\mathbb{R}$ to denote an admissible heuristic function, similar to standard A*. Lines 2–6 initialize the variables, where the function $p(v,\ell)$ returns the ancestor of node $v$ in layer $\ell$. Lines 7–24 execute COA*, which will be detailed in Section IV-B. Notably, Line 14 computes the semantic class of the edge $(v,u)$ based on the semantic classes of nodes $v$ and $u$. If $\ell=0$, the nodes’ semantics are determined from perception data, whereas for $\ell\neq 0$, they are predicted using the methods described in Section V. Additionally, $G^{0^{\prime}}(u)$ is the induced subgraph of node $u$ in layer $\ell=0$ given by $G^{0^{\prime}}(u)=\{u^{\prime}\in V^{0}:p(u^{\prime},\ell)=u\}$. Finally, Lines 25–27 refine the path by pruning nodes that do not share an ancestor with the paths in the higher layers.
@@ IV-B Class-ordered A*
In this section, we present an overview of the simplified COA* implementation used in HCOA* (Lines 7–24 in Algorithm 1) to help the reader understand the simplifications made on the baseline algorithm. Further details, along with proofs of the COA* optimality and completeness, can be found in [9].
Consider a single-layer, weighted, semantic graph $G^{\ell}$. In this algorithm, each node $v$ is characterized by the triple $(q,g,\theta)$, where $q$ is the predecessor node, $g:V^{\ell}\to\mathbb{R}^{+}$ is the cost-to-come function, and $\theta_{V}:V\to\mathbb{N}^{K}$ specifies the number of edges of each semantic class along the path up to node $v$. We extend $\theta_{V}$ to single edges by defining $\theta_{E}:E^{\ell}\to\mathbb{N}^{K}$ as an one-hot vector with all elements zero except at $\phi_{E}$. The subscript is omitted in Algorithm 1. The function PopNode selects the next node to expand lexicographically, and Path constructs the optimal path by backtracking through parent nodes from the goal. Figure 2 presents the path of COA* in a grid world environment.
Figure 2: (a) Grid environment with three semantic classes: road, grass, river. S denotes the starting cell while G denotes the target; (b) Associated graph of the environment and optimal path produced by Class-ordered A*.
Figure 3: Proposed GNN architecture for node $P\in V^{\ell}$ ($\ell\neq 0$) classification, utilizing the semantic classes (green, blue, red) of nodes in $\ell=0$. The dataset $\mathcal{D}=\{G^{0^{\prime}}(P),\phi_{V}^{\ell}(P)\}$ is generated by running COA* on subgraphs of induced nodes. The network consists of two MLPs for pre-processing and post-processing, three GNN layers with skip connections, and an average pooling operator. Training is conducted using cross-entropy loss over the semantic classes.
@@ IV-C Performance Guarantees
The following propositions establish the algorithm’s completeness along with sufficient conditions for its optimality. For both propositions, we assume that the assumptions for completeness and optimality of COA* hold [9]. Proposition 2 relies on two key assumptions. While these assumptions may seem restrictive, they often hold in indoor environments, as discussed in Section VI.
Proposition 1
@@
Proof:
Let $\pi^{0}$ exists. The structure of $G$ ensures that for each layer $\ell\in L$, there exists a corresponding path $\pi^{\ell}=\pi(v_{s}^{\ell},v_{g}^{\ell})$, where $v_{s}^{\ell}=p(v_{s},\ell)$ and $v_{g}^{\ell}=p(v_{g},\ell)$.
At each subgraph $G^{\ell}$, COA* is guaranteed to find the optimal path $\pi^{\ell*}$, if one exists [9]. Suppose, ad absurdum, that the graph pruning performed in Lines 25–27 results in $\Pi(v_{s}^{\ell-1},v_{g}^{\ell-1})=\varnothing$, preventing COA* from finding a solution in layer $\ell-1$. This would imply that $\pi^{\ell*}$ is disconnected, contradicting the completeness of COA*. ∎
Proposition 2
@@
Proof:
Define $G^{0^{\prime}}\subseteq G^{0}$ as the pruned graph after $n$ iterations of HCOA*. Suppose, ad absurdum, that a segment of the optimal path $\pi^{*}(v_{s},v_{g})$ was removed, i.e., there exists a sub-path $\widehat{\pi}\subseteq\pi^{*}(v_{s},v_{g})$ such that $\widehat{\pi}\not\subset G^{0^{\prime}}$. Two cases arise: (i) The parents of $\widehat{\pi}$ in layer $\ell=1$ form a cycle that starts and ends at a parent of a node in $\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}$. This contradicts the second assumption, which states that the nodes of the optimal path at $\ell=0$ where the starting and goal node share the same parent, must all have the same parent. (ii) The parents of $\widehat{\pi}$ in layer $\ell=1$ are part of a sub-path that starts at the parent of one node in $\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}$ and ends at a different parent of another node in $\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}$. This contradicts the first assumption, as it implies the existence of two distinct paths in layer $\ell=1$, violating uniqueness. ∎

### v2

@@ Abstract
This paper addresses the problem of robot navigation in mixed geometric/semantic 3D environments. Given a hierarchical representation of the environment, the objective is to navigate from a start position to a goal, while satisfying task-specific safety constraints and minimizing computational cost. We introduce Hierarchical Class-ordered A* (HCOA*), an algorithm that leverages the environment’s hierarchy for efficient and safe path-planning in mixed geometric/semantic graphs. We use a total order over the semantic classes and prove theoretical performance guarantees for the algorithm. We propose three approaches for higher-layer node classification based on the semantics of the lowest layer: a Graph Neural Network method, a k-Nearest Neighbors method, and a Majority-Class method. We evaluate HCOA* in simulations on two 3D Scene Graphs, comparing it to the state-of-the-art and assessing the performance of each classification approach. Results show that HCOA* reduces the computational time of navigation by up to 50%, while maintaining near-optimal performance across a wide range of scenarios.
Index Terms:
Autonomous Vehicle Navigation, Motion and Path Planning, AI-Enabled Robotics.
@@ I Introduction
As robotic sensing technologies advance, enabling robots to perceive vast and diverse information, two fundamental questions arise: What information from this extensive data stream is most important for a given task? and, second, how can the robot effectively utilize this information for decision-making? Hierarchical semantic environment representations, such as 3D Scene Graphs (3DSGs) [1, 2, 3], provide rich and structured abstractions that mirror human-like reasoning, thus facilitating the selection and organization of information.
@@
Previous research in hierarchical path-planning has primarily addressed the first question [4, 5, 6, 7]. In [4] the authors introduce Hierarchical Path-Finding A*, a hierarchical A* [8] variant for grid-based maps. Their approach partitions the map into clusters with designated entrance points, which are used for high-level path-planning. The authors in [5] propose a hierarchical graph search algorithm for graphs with edge weights represented as intervals. In [7], the authors introduce S-Nav, a hierarchical path-planning algorithm for navigation in Situational Graphs, a multi-layer graph representation of indoor environments built on top of 3DSGs.
Semantic path-planning has focused on the second question above by incorporating semantics into the decision-making process. In [9], the authors propose SMaNa, a semantic-aware mapping and navigation framework that employs A* with edge weights computed using node distances and semantic labels. To avoid reliance on manually crafted, user-defined cost functions, several works adopted the concept of total order over semantic classes. In [10], a weighted function combining edge cost and semantic class is introduced to determine optimal paths while respecting a total semantic order. However, this approach requires global graph properties, making it computationally expensive. To address this limitation, the authors in [11] propose Class-ordered A* (COA*), an extension of traditional A* that incorporates semantic information, based on a total order. The algorithm is shown to be both complete and optimal under the defined total order and is validated in various planning scenarios. This work was later extended to dynamic environments in [12], where LPA* [13] was adapted for online replanning in dynamic maps with semantic information.
Building on the ideas introduced in [5] and [11], this paper proposes a hierarchical class-ordered path-planning algorithm that uses a total order on semantic classes. The notion of total order is applicable to a wide range of scenarios, including navigation in partially known environments [11] and safety-critical applications. In this work, we focus on encoding a set of safety constraints as relationships between semantic categories, which are implicitly enforced during path-planning.
@@
Navigation within hierarchical semantic environments has also been explored in prior work. In [15], the authors present IntelliMove, a hierarchical semantic map framework along with a semantic path-planner. The edge weights in the graph encode spatial relationships, such as distances between rooms or objects, and Dijkstra’s algorithm is used to compute optimal paths. Furthermore, [16] addresses Task and Motion Planning in 3DSGs using a three-level hierarchical planner consisting of a discrete task planner, a navigation planner for path-planning, and a low-level planner. Their approach focuses on optimizing task planning to reduce computational cost, while path-planning operates on the unpruned 3DSG.
@@
Main Contributions: Our approach integrates task semantics directly into the path-planning process within a unified algorithm. By leveraging the environment’s hierarchy, we reduce computational resource demands while ensuring efficient and safe navigation. Our key contributions are as follows:
@@
    We propose three methods for node classification on the higher layers of the environment’s hierarchy: a Majority-Class (MC) method, a k-Nearest Neighbors (kNN) method, and a GNN method.
@@
    We validate our approach on two publicly available 3DSG datasets, demonstrating its effectiveness in real-world scenarios. [^1] The code is available at https://github.com/epsomiadis3/HCOA-star.
@@ II Preliminaries
We model the environment as a hierarchical semantic graph (HSG). Specifically, let $G=(V,E,\mathcal{K})$ be a graph with $n$ layers, where $V$ is the set of nodes, $E\subseteq V\times V$ is the set of edges, and $\mathcal{K}={1,\ldots K}$ represents a set of semantic classes ordered in decreasing priority. This priority is characterized by task-specific requirements (e.g., avoid certain areas or objects, or prefer known over unknown regions). We denote the set of layers as $L$, where $\ell=0$ is the lowest abstraction (e.g., sensor) layer and $\ell=n$ is the highest abstraction layer (root). Each layer $\ell$ forms a connected weighted subgraph $G^{\ell}=(V^{\ell},E^{\ell},\mathcal{K})\subset G$, with an associated weight function $w^{\ell}:E^{\ell}\to\mathbb{R}^{+}$, which corresponds to Euclidean distance in this work. We assume that each node in layer $\ell-1$, for $\ell=1,\ldots n$, is connected to a single node in layer $\ell$, which we refer to as its parent node. More generally, we define an ancestor as a node’s parent or any higher-layer predecessor in the hierarchy. We define the projection function $p:V\times L\to V$ that maps each node to its corresponding ancestor in layer $\ell$.
Each node in $\ell=0$ is assigned a semantic class based on the given perception data and the specific task using the function $\phi_{V}^{0}:V^{0}\to\mathcal{K}$. We extend the labeling function $\phi_{V}^{0}$ to layers $\ell\neq 0$. Details on the computation of $\phi_{V}^{\ell}$ for $\ell\neq 0$ are given in Section V. Additionally, we define the edge classification function $\phi_{E}^{\ell}:E^{\ell}\to\mathcal{K}$ by $\phi_{E}^{\ell}(e)=\max(\phi_{V}^{\ell}(v),\phi_{V}^{\ell}(u))$ for $e=(v,u)$, thereby overestimating the edge class.
The environment model reflects the structure of 3DSGs, as shown in Figure 1. For an overview of the advantages of this perceptual model and its construction, we refer the interested reader to [2]. In our framework, in addition to the spatial layers, we also utilize the object layer of the 3DSG to encode safety constraints and to infer the nodes’ semantic classes, enabling more context-aware and task-relevant navigation.
Let $\Pi(v_{s}^{\ell},v_{g}^{\ell})$ denote the set of all acyclic paths in $G^{\ell}$ from $v_{s}^{\ell}\in V^{\ell}$ to $v_{g}^{\ell}\in V^{\ell}$, and let $\Pi_{k}$ be the subset of all paths in $\Pi$ for which the least favorable (i.e., highest) edge class is exactly $k$. Formally,
$\Pi_{k}=\{\pi\in\Pi:N(\pi,k)>0;\>N(\pi,k^{\prime})=0,\forall k^{\prime}>k\},$ (1)
where $N(\pi,k)=|\{e\in\pi:\phi_{E}^{\ell}(e)=k\}|$ is the number of edges of class $k$ in path $\pi$. We further define $\Pi_{k}^{i}\subseteq\Pi_{k}$ consisting of all paths in $\Pi_{k}$ that contain exactly $i$ edges of class $k$, that is $\Pi_{k}^{i}=\{\pi\in\Pi_{k}:N(\pi,k)=i\}$. To enable a consistent comparison of paths across different classes, we impose a total order such that $k<\ell\Rightarrow\Pi_{k}^{i}\prec\Pi_{\ell}^{j}$ $\text{for all }i,j$ and $i<j\Rightarrow\Pi_{k}^{i}\prec\Pi_{k}^{j}$. This order ensures that any two paths with the same start and goal nodes can be compared.
@@
4:  $g(v^{\ell})\leftarrow\infty;\hskip 5.0pt\theta(v^{\ell})\leftarrow\infty\cdot\mathbf{1}_{K},\quad\forall v^{\ell}\in G^{\ell^{\prime}}\setminus\{v_{s}^{\ell}\}$
5:  PredecessorMap $\leftarrow\varnothing$ $\triangleright$ Track path reconstruction
6:  $Q\leftarrow\{v_{s}^{\ell}\}$ $\triangleright$ Initialize priority queue
@@
26:  for all $\hat{\ell}$$\ell$ do
27:   $G^{\hat{\ell}^{\prime}}\leftarrow\{v\in G^{\hat{\ell}^{\prime}}:p(v,\ell)\in\pi^{\ell}\}$
@@ III Problem Formulation
Consider a mobile robot tasked with navigating a complex 3D environment where certain regions have lower traversal priority. The robot is provided with a HSG of the environment (e.g., 3DSG), as outlined in Section II, where the semantic classes of the nodes in layer $\ell=0$ are assigned based on perception, and tailored to the specific task (e.g., safety constraints). Let $v_{s}\in V^{0}$ and $v_{g}\in V^{0}$ denote the robot’s starting and goal nodes, respectively. The objective is to determine the shortest path in $\ell=0$ while minimizing traversal through the least favorable edges. Formally,
$\pi^{*}(v_{s},v_{g})=\operatorname*{arg\,min}\limits_{\pi\in\Pi^{*}(v_{s},v_{g})}{\sum\limits_{e\in\pi}w^{0}(e)},$ (2a)
@@ III-A Problem Statement
We propose a hierarchical semantic path-planning algorithm for efficient robot navigation in large-scale environments. The algorithm operates top-down across the layers of the HSG, iteratively computing the optimal semantic path at each layer while pruning nodes that are not included in the path. Additionally, we introduce three methods for node classification to predict the semantic classes of higher-layer nodes: a Majority-Class, a kNN, and a GNN method.
@@ IV-A Hierarchical Class-ordered A*
We introduce Hierarchical Class-ordered A* (HCOA*), a hierarchical semantic path-planning algorithm for HSGs. The algorithm initiates planning at layer $\ell=n-1$, and recursively proceeds to lower layers. At each layer, it prunes the HSG based on the computed path from the previous layer and applies the same planning algorithm to the pruned HSG. Within each layer, the algorithm utilizes Class-ordered A* (COA*) [11], which finds the shortest path while minimizing the number of least favorable edges through lexicographic comparison. By leveraging the hierarchy, the algorithm performs multiple searches on smaller subgraphs rather than a potentially computationally expensive search over $G^{0}$.
HCOA* is presented in Algorithm 1. We use the function $h:V\to\mathbb{R}$ to denote an admissible heuristic function, similar to standard A*. Lines 2–6 initialize the variables, where the function $p(v,\ell)$ returns the ancestor of node $v$ in layer $\ell$. Lines 7–24 execute COA*, which will be detailed in Section IV-B. Notably, Line 14 computes the semantic class of the edge $(v,u)$ based on the semantic classes of nodes $v$ and $u$. If $\ell=0$, the nodes’ semantics are determined from perception data and the task, whereas for $\ell\neq 0$, they are predicted using the methods described in Section V. Additionally, $G^{0^{\prime}}(u)$ is the induced subgraph of node $u\in V^{\ell}$ on layer 0 given by $G^{0^{\prime}}(u)=\{u^{\prime}\in V^{0}:p(u^{\prime},\ell)=u,\>\text{where}\>u\in V^{\ell}\}$, while $G^{\ell^{\prime}}$ is the pruned graph at layer $\ell$ (Line 27). Lines 25–27 refine the HSG by pruning nodes that do not share an ancestor with the paths in the higher layers.
@@ IV-B Class-ordered A*
In this section, we present an overview of the simplified COA* implementation used in HCOA* to help the reader understand the simplifications made to the baseline algorithm. Further details, along with proofs of the COA* optimality and completeness, can be found in [11].
Consider a single-layer, weighted, semantic graph $G^{\ell}$. We characterize each node $v$ by the triple $(q,g,\theta_{V})$, where $q$ is the predecessor node (saved in the PredecessorMap in Algorithm 1), $g:V^{\ell}\to\mathbb{R}^{+}$ is the cost-to-come function, and $\theta_{V}:V^{\ell}\to\mathbb{N}^{K}$ specifies the number of edges of each semantic class along the path up to node $v$. We extend $\theta_{V}$ to single edges by defining $\theta_{E}:E^{\ell}\to\mathbb{N}^{K}$ as an one-hot vector with all elements zero except at $\phi_{E}$. The subscript is omitted in Algorithm 1. The function PopNode selects the next node to expand, and Path constructs the optimal path by backtracking through the predecessor nodes from the goal.
Figure 2: (a) Grid world with three semantic classes: road, grass, river. S is the starting cell while G is the target; (b) Graph of the environment and path produced by COA*.
Figure 2 presents an example of COA* in a grid world, where the robot must navigate from $\mathbf{S}$ to $\mathbf{G}$ while satisfying safety constraints prioritized in descending order: (i) avoid water, and (ii) avoid grass. To enforce these constraints, we assign semantic labels to nodes via $\phi_{V}^{\ell}$.
@@ IV-C Performance Guarantees
The following propositions establish the algorithm’s completeness along with sufficient conditions for optimality. For both propositions, the assumptions for completeness and optimality of COA* hold [11]. We conclude the section with a discussion on the algorithm’s computational complexity.
Proposition 1.
@@
Proof.
Let $\pi^{0}$ exist. The structure of $G$ ensures that for each layer $\ell\in L$, there exists a corresponding path $\pi^{\ell}=\pi(v_{s}^{\ell},v_{g}^{\ell})$, where $v_{s}^{\ell}=p(v_{s},\ell)$ and $v_{g}^{\ell}=p(v_{g},\ell)$. At each subgraph $G^{\ell}$, COA* is guaranteed to find the optimal path $\pi^{\ell*}$, if one exists [11]. Suppose, ad absurdum, that the graph pruning performed in Lines 25–27 results in $\Pi(v_{s}^{\ell-1},v_{g}^{\ell-1})=\varnothing$, preventing COA* from finding a solution in layer $\ell-1$. This implies that $\pi^{\ell*}$ is disconnected, contradicting COA*’s completeness. ∎
Proposition 2.
@@
Proof.
Define $G^{0^{\prime}}\subseteq G^{0}$ as the pruned graph after $n$ iterations of HCOA*. Suppose, ad absurdum, that a segment of the optimal path $\pi^{*}(v_{s},v_{g})$ has been removed, i.e., there exists a sub-path $\widehat{\pi}\subseteq\pi^{*}(v_{s},v_{g})$ such that $\widehat{\pi}\not\subset G^{0^{\prime}}$. Two cases arise: (i) The parents of $\widehat{\pi}$ in layer $\ell=1$ form a cycle that starts and ends at a parent of a node in $\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}$. This contradicts the second assumption, which states that the nodes of the optimal path at $\ell=0$, where the starting and goal node share the same parent, must all have the same parent. (ii) The parents of $\widehat{\pi}$ in layer $\ell=1$ are part of a sub-path that starts at the parent of one node in $\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}$ and ends at a different parent of another node in $\pi^{*}(v_{s},v_{g})\cap G^{0^{\prime}}$. This contradicts the first assumption, as it implies the existence of two distinct paths in layer $\ell=1$, violating uniqueness. ∎
The two assumptions of Proposition 2 may seem restrictive at first, but they often hold in indoor environments. Figure 3 shows an example, where $\mathbf{G}$ denotes the goal, $R(\cdot)$ denotes a room, and $C_{1}$, $C_{2}$ represent specific conditions, namely $C_{1}$: open door and $C_{2}$: dangerous area. When both $C_{1}$ and $C_{2}$ are false, Proposition 2 holds, and HCOA* successfully finds the optimal path (shown in green). However, when $C_{1}$ is true, the path produced by HCOA* (still shown in green) becomes suboptimal. This occurs because the room is represented, in this example, by its center node. For large rooms such as corridors (e.g., $R(1)$), this representation becomes misleading. Additionally, if $C_{2}$ is true, the optimal path requires moving from $R(2)$ to $R(1)$ and then back to $R(2)$. This cycle cannot be captured by COA* operating at the room layer, as it computes acyclic paths. Hence, HCOA* produces a suboptimal path in the places layer in this case (shown in blue).
In general, the spatial representation of nodes in higher layers, as well as their computed semantic classes, can significantly affect HCOA*’s performance, potentially leading to suboptimal solutions. Nevertheless, as demonstrated in Section VI, HCOA* manages to find the optimal path in most tested scenarios. One potential improvement is to refine the representation of rooms by associating multiple nodes with each room, for example, using nodes near doors rather than using a single node at the geometric center of the room.
Figure 3: Example demonstrating Proposition 2. $\mathbf{G}$ denotes the goal location, $R(\cdot)$ indicates rooms, and $C_{1}$, $C_{2}$ represent specific conditions: $C_{1}$: open door, and $C_{2}$: dangerous area.
Figure 4: Proposed GNN architecture for node $P\in V^{\ell}$ ($\ell\neq 0$) classification, utilizing the semantic classes (green, blue, red) of nodes in $\ell=0$. The dataset $\mathcal{D}=\{G^{0^{\prime}}(P),\phi_{V}^{\ell}(P)\}$ is generated by running COA* on induced subgraphs of nodes. The network consists of two MLPs for pre-processing and post-processing, three GNN layers with skip connections, and an average pooling operator. Training is conducted using cross-entropy loss over the semantic classes.
Regarding the computational complexity of the algorithm, note that, in the worst case, COA* with an admissible heuristic has the same complexity as Dijkstra’s algorithm, which is $\mathcal{O}(|E^{0}|+|V^{0}|\log|V^{0}|)$ when implemented with a Fibonacci heap. Hence, the worst-case complexity of HCOA* is given by $\mathcal{O}\left(\sum_{\ell=0}^{L}\left(|E^{\ell^{\prime}}|+|V^{\ell^{\prime}}|\log|V^{\ell^{\prime}}|\right)\right)$, where $|V^{\ell}{}^{\prime}|$ and $|E^{\ell}{}^{\prime}|$ denote the number of vertices and edges, respectively, of the pruned graph at each layer $\ell$.